OpenStudy (thedj4jc):

Find the distance between the points (-4, -5) and (3, -1).

OpenStudy (butterfield1215):

do you know what the formula to find this is

OpenStudy (thedj4jc):

\[\sqrt{65}\] \[\sqrt{37}\] \[\sqrt{17}\]

OpenStudy (thedj4jc):

No

OpenStudy (anonymous):

@thedj4JC make it into a Graph and That Would Help While Using the pythagorean theorem

OpenStudy (butterfield1215):

ok

OpenStudy (♪chibiterasu):

\[Distance = \sqrt{x_2 - x_1}^2 + \sqrt{y_2 - y_1}^2\]

OpenStudy (♪chibiterasu):

You don't need to find the square root of the result, you can leave the radical in since the options don't seem to do that.

OpenStudy (thedj4jc):

I can't get the square roots to work with negatives, why is that?

OpenStudy (thedj4jc):

@♪Chibiterasu ??